.Viable plate count

Using the formula: cfu/ml = (no. of colonies x dilution factor) / volume of culture plate In plate 10-4 there were 150 colonies counted.

Experimental process: Pipet 0.25 mL of the S. cerevisiae solution into the test tube labeled 10-1. Pipet the solution up and down several times to ensure all organisms are rinsed from the pipet into the test tube.

Pipet 0.25 mL of the 10-1 S. cerevisiae solution into the test tube labeled 10-2.

Pipet the solution up and down several times to ensure all organisms are rinsed from the pipet into the test tube. Repeat the transfer process to pipet the S. cerevisiae solution from the 10-2 tube to the 10-3 tube; from the 10-3 tube to the 10-4 tube; from the 10-4 tube to the 10-5 tube; and from the 10-5 tube to the 10-6 tube.

The yeast was added to 2.25ml of distilled water. The a pipet 0.125 mL (4 drops) of each solution was added to the corresponding plate.

Calculate:

- Dilution Plate CFUs

- Dilution Factor of Viable Plate

- Volume of Diluent added to Plate

- CFU/mL of Original Sample:

- CFU/(Volume x Dilution Factor)
Answer #1

1 colonies on 10-4 dilution = 150 (given)

2 dilution factor of viable plate - 10-4

3 volume of diluent added to plate:- 0.25 ml

4 CFU/ml of original sample =

Since the volume taken was 0.25 ml therefore-

colonies on 10-4 dilution = 150 colonies/0.25 ml = 600 cfu/ml

so no. of microbes in original solution = 600/10-4 =600x 104 cfu/2.25ml = 266.6 x 104 cfu/ml (original sample was prepared in 2.25 ml)

5 CFU/(Volume x Dilution Factor)= 150/(0.25 x 10-4)

= 6 x 106 cfu/ml